OpenStudy (anonymous):

uh thats not right. its: 4x + (-4) =-36

OpenStudy (anonymous):

x=-8

OpenStudy (anonymous):

ooh thanks!

OpenStudy (anonymous):

4x-4=-36 4x=-32 x=-32/4=-8